OpenStudy (anonymous):

Solve: 4 + 2(x − 1) = 5x − (x + 10)

OpenStudy (anonymous):

Distribute 2 to (x-1) first on left hand side...

OpenStudy (anonymous):

wat next?

OpenStudy (anonymous):

Do this much first and show me what did you get?

OpenStudy (anonymous):

4 + 2(x − 1) = 5x − (x + 10) 4+ 2x-2= 5x-x-10 2+2x= 4x-10 - 2x -2x 2=2x-10 +10 +10 12=2x /2 /2 6=x